OpenStudy (anonymous):

convert 764 base 8 numbers to base 2

OpenStudy (shubhamsrg):

Try to convert to base 10 first

OpenStudy (agent0smith):

\[764_{8} = 7 \times 8^2 + 6 \times 8 +4 = 500\]

OpenStudy (agent0smith):

Find the biggest power of 2 that goes into 500 first...